OpenStudy (ttop0816):

am i correct?? sin^2 60° + cos^2 30° = _____ 1/2 3/4 1 1 1/2 <

OpenStudy (anonymous):

sin 60 = cos 30 = \[\frac{ \sqrt{3} }{ 2 } \]

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (ttop0816):

no ): mind helping me out?

OpenStudy (anonymous):

OK \[{(\frac{ \sqrt{3} }{ 2 })}^2=\frac{ \sqrt{3}^2 }{ 2^2 }=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

sin^2 60° + cos^2 30° =3/4 +3/4=6/4=3/2

OpenStudy (ttop0816):

ohhh then the answer would equal to 3/4!

OpenStudy (anonymous):

is 3/4=3/2 ?

OpenStudy (ttop0816):

ohhh sorry nvm 1 1/2!

OpenStudy (anonymous):

Yes that is it, I hate that notation. It always confused me. However it is true 1 1/2=3/2

OpenStudy (ttop0816):

thank you!! @Andras oh and do you know if this is correct also?? Convert 28.7251° to degrees, minutes, and seconds. Round to the nearest second. i got 28 degrees 43' 30.36''?? and how would i round it to the nearest second??

OpenStudy (anonymous):

Here is your cheat: http://www.satsig.net/degrees-minutes-seconds-calculator.htm :)

OpenStudy (anonymous):

\[\sin ^2(60 {}^{\circ})+\cos ^2(30 {}^{\circ}) \]\[\left(\frac{\sqrt{3}}{2}\right)^2+\left(\frac{\sqrt{3}}{2}\right)^2 \]\[\frac{3}{4}+\frac{3}{4}=\frac{3}{2} \]